OpenStudy (linaminas):

Is this right ? Use the distributI've property to rewrite the expression 4(7+3) I got 31

OpenStudy (mariamikayla):

4(7+3)= 4*7+4*3= =28+12= =40

OpenStudy (linaminas):

Oh okay you a have to mult. 4 by both , thank you!

OpenStudy (mariamikayla):

yes, that the property of distribution :)) you're welcome :D

OpenStudy (linaminas):

Thank you I'm rusty on alg. ): haha

OpenStudy (mariamikayla):

you just have to practice a little bit more and you'll understant algebra better ;) :D knowing maths come after practicing :))